OpenStudy (anonymous):

Use the point-slope form to write an equation of the line with the given properties. Leave the answer in point-slope form. Slope −7; passes through (9, −2)

OpenStudy (anonymous):

I got y=-7x+61 but it says its wrong

Nnesha (nnesha):

point slope form is \[\huge\rm y-y_1=m(x-x_1)\] where is m is slope and \[\large\rm (x_1,y_1)\] is the the order pair (9,-2) so plug in!

Nnesha (nnesha):

that's the slope intercept form but you have to write in `point-slope` form `Leave the answer in point-slope form`

OpenStudy (anonymous):

oh okay thankyou

Nnesha (nnesha):

np :=) let me know if you want me to check ur answer :=)

OpenStudy (anonymous):

How do I put it into point-slope form? Do I just put the answer in it?

Nnesha (nnesha):

this is the point slope form \[\huge\rm y-y_1=m(x-x_1)\] where is m is slope and \[\large\rm (x_1,y_1)\] is the the order pair (9,-2) so substitute m for -7 and (x_1 ,y_1) for (9,-2)